Sn = 8 + 10 + 12 + 14

1) Express T2 in terms of s1 and s2
2) express T3 in terms of S3 and S2
3)Express T4 in terms of S3 and S4

Answer by ikleyn(52788)

.
Sn = 8 + 10 + 12 + 14
1) Express T2 in terms of s1 and s2
2) express T3 in terms of S3 and S2
3)Express T4 in terms of S3 and S4
~~~~~~~~~~~~~~~~~~~

1)  T2 = S2 - S1



2)  T3 = S3 - S2



3)  T4 = S4 - S3

Solved.     //     All questions are answered.